If the total positive charge is Q = 1.62×10−6 C, what is the magnitude of the...

uploaded imageIf the total positive charge is Q = 1.62×10−6 C, what is the magnitude of the electric field caused by this charge at point P, a distance d = 1.53 m from the charge?
Answer #1

electric field E =k (Q/d^2)   .......... (1)

where k =8.99*10^9 N.m^2/C^2

charge Q =1.62*10^-6 C

distance d =1.53 m

substitute these values in eq (1), we get

electric field E =(8.99*10^9)[(1.62*10^-6)/(1.53)^2]

E =6.22*10^3 N/C

E =6.22 kN/C